OpenStudy (anonymous):

I have a text book which writes a function of time can be written as (V0 is the frequency) F(t)=A02+∑m=1∞Rmcos(2πmv0t+ϕm) and with Am=Rmcosϕm Bm=Rmsinϕm then Function can be rewrite as: F(t)=A02+∑m=1∞e2πmv0t(Am−iBm) It seems there are some internal relations between these two functions and I have tried expand and elaborate these two functions but can't work out a direct proof can someone help me out?

hartnn (hartnn):

*

OpenStudy (anonymous):

I have a text book which writes a function of time can be written as (V0 is the frequency) \[F(t)=\frac{ A _{0} }{ 2 }+\sum_{m=1}^{\infty} R _{m} \cos(2 \pi mv _{0}t+\phi _{m})\] and with \[A_{m}=R_{m}\cos \phi_{m} \] \[B_{m}=R_{m}\sin \phi_{m} \] then Function can be rewrite as: \[F(t)=\frac{ A_{0} }{2 }+\sum_{m=1}^{\infty}e ^{2\pi mv_{0}t}(A_{m}-iB_{m})\] It seems there are some internal relations between these two functions and I have tried expand and elaborate these two functions but can work out a direct proof can someone help me out?

hartnn (hartnn):

i haven't tried it on paper, but these relation should lead you to what u need, \(\cos (A+B) = \cos A\cos B-\sin A \sin B , \quad \cos (2\pi m v_ot+\phi_o)=...?\) \(\cos x = (e^x+e^{-x})/2 \\ \sin x =(e^x-e^{-x})/2i \) try it and ask if you get stuck...

OpenStudy (anonymous):

@hartnn close but those should be imaginary exponentials @shadow_wxh This 'relationship' is merely Euler's:$$e^{i\theta}=\cos\theta+i\sin\theta\\\cos\theta=\frac12\left(e^{i\theta}+e^{-i\theta}\right)\\\sin\theta=\frac1{2i}\left(e^{i\theta}-e^{-i\theta}\right)$$

hartnn (hartnn):

yes, i meant that :)

OpenStudy (anonymous):

To see their first step,$$\cos(2\pi mv_0t+\phi_m)=\cos(2\pi mv_0 t)\cos(\phi_m)+\sin(2\pi mv_0 t)\sin\phi_m$$

hartnn (hartnn):

oldrin, let shadow do the work!

OpenStudy (anonymous):

$$e^{2\pi mv_0 t}(A-iB_m)=R_m(\cos(2\pi mv_0 t)+i\sin(2\pi mv_0 t))(\cos\phi_m-i\sin\phi_m)=\cdots?$$

OpenStudy (anonymous):

I have done the above but missing any further lead

OpenStudy (anonymous):

\[R_{m}\cos(2pimv_{0}t+\phi_{m})=R_{m}\cos \phi_{m}\cos(2 \pi mv_{0}t)-R_{m}\sin \phi_{m}\sin(2 \pi mv_{0}t)\]

OpenStudy (anonymous):

\[=A_{m}\cos(2 \pi mv_{0}t)-B_{m}\sin(2 \pi mv_{0}t)\]

OpenStudy (anonymous):

Good so far...

OpenStudy (anonymous):

hmm gonna need a theory there

OpenStudy (anonymous):

Any body give me a hint?

OpenStudy (anonymous):

Ok I try work backwards\[e ^{2 \pi m v_{0}t}(A_{m}-iB_{m})=[\cos(2 \pi mv_{0}t+i \sin(2\pi mv_{0}t)](A_{m}-iB{m})\]\[=A_{m}\cos(2\pi mv_{0}t)+iA_{m}\sin(2\pi mv_{0}t)-iB_{m}\cos(2\pi mv_{0}t)+B_{m}\sin(2\pi mv_{0}t)\] if assume the question is true then \[e ^{2 \pi m v_{0}t}(A_{m}-iB_{m})=A_{m}\cos(2\pi mv_{0}t)+B_{m}\sin(2\pi mv_{0}t)\] which left \[iA_{m}\sin(2\pi mv_{0}t)-iB_{m}\cos(2\pi mv_{0}t)=0\] \[iR_{m} cos\phi_{m}sin(2 \pi mv_{0}t)-iR_{m}sin\phi _{m}cos(2\pi mv_{0}t)=0\] \[iR_{m}sin(2\pi mv_{0}t-\phi_ {m})=0\] \[2\pi mv_{0}t=\phi_{m}\] ?????Really?

OpenStudy (anonymous):

Corrrection there is i missing in the above \[e^{2\pi imv_{0}t}\]

OpenStudy (anonymous):

come on people try this out

OpenStudy (anonymous):

Tried many way coudln't prove it unless m goes from minus infinity to positive infinity the closest equation is \[F(t)=\sum_{m=-\infty}^{+\infty}C_{m}e^{2\pi imv_{0}t}\] where \[C_{m}=\frac{A_{m}-iB_{m}}{2}\] But that doesnt make them equal so i'll say there is some mistake i the text book.

OpenStudy (anonymous):

OK anyone wants to challenge this or i'll just close it

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

You have any better suggestion?